# DreamVLA: A Vision-Language-Action Model Dreamed with Comprehensive World Knowledge

### v2

@@ 2.1 Vision–Language–Action Models
The earliest VLA [16, 71, 2, 72, 73, 74] lay the foundation by combining pretrained vision-language representations with task-conditioned policies for manipulation and control. Inspired by the recent advances of Large Language Models [75, 76, 77, 78] and multimodal large language models [28, 26, 79, 65, 80] and the emergence of large-scale robot datasets [12, 81, 82, 83], VLA has become a trend in robot learning. RT series [2, 84, 85] is the pioneer attempt to fine-tune the MLLM on robot demonstration datasets, resulting in strong accuracy and generalization. Building on this foundation, many advanced techniques [30, 32, 13, 1, 33, 34, 73, 35, 36, 37, 86, 87, 88, 38, 89] are developed to boost the performance. Meanwhile, considering the advantage of the diffusion model in modeling multi-peak, some researchers [90, 91, 92, 93, 94] employ different architectures to sample action from noise conditioned on observation, task instruction, and robot prior knowledge. Given on this manner which directly maps observation and instruction to action lacks reasoning steps like LLM [62], most existing methods [43, 5, 44, 45, 46, 47, 48, 49] leverage a copilot image/video generation model to generate future frames then predict action sequences conditioned on goal images. However, the above methods still need an extra generation model, which introduces inference time and computation load. Therefore, several methods [56, 57, 58, 59, 60, 61] integrate pixel-level image forecasting with the action prediction in a single framework, which exploits the synergy of prediction and planning. Despite success, these methods naturally exhibit limitations in redundant reconstruction [95], and suffer from a lack of spatial and semantic information.
@@ 2.2 Knowledge Forecasting for Robotics
Learning future world knowledge for robot training has increasingly become popular to enable policies for achieving an action-forecasting loop. Early attempts [49, 19, 14, 43, 51, 50, 96] to implement this based on off-the-shelf world models [90, 97, 53] and feed the forecasting of images or states into the robot policy model to conduct inverse dynamics. This two-stage training strategy is easy to implement but is limited by the performance of world models. More advanced solutions [58, 56, 88, 98] incorporate forecasting and action planning in an end-to-end manner, requiring the policy to predict actions along with future states. This line of work demonstrates better performance and generalization. However, the future states are limited to redundant visual information [63, 64, 99, 69, 100, 66] or monotonous states [21, 48]. In contrast to previous work, DreamVLA proposes to predict future knowledge in an efficient (dynamic region) and effective (comprehensive knowledge) way, demonstrating strong performance and generalization.
@@ 3.1 Problem Definition and Notation
[…]
@@
[…]
@@ 3.3 Comprehensive World Knowledge Prediction
$\sum_{(z_{i},\tilde{z}_{i})\in\mathcal{D}}\log\mathrm{P}(x_{i}|\tilde{x}_{i})\geq\sum_{(x_{i},\tilde{x}_{i})\in\mathcal{D}}\Big{(}\mathbb{E}_{z_{i}\sim\mathrm{Q}_{\phi}(\mathbf{z}|x_{i})}\big{[}\log\mathrm{P}_{\psi}(x_{i}|z_{i})\big{]}-D_{\mathrm{KL}}\big{[}z,\mathrm{P}_{\theta}(\mathbf{z}|\hat{z_{i}})\big{]}\Big{)},$ (4)
where $\mathrm{P}_{\psi}(x|z)$ is the tokenizer decoder to recover origin data, $\hat{z_{i}}=\mathrm{Q}_{\phi}(\mathbf{z}|\tilde{x}_{i})$ denotes the masked motion region tokens from masked data and $\mathrm{P}_{\theta}(z|\hat{z_{i}})$ reconstructs masked tokens in an autoencoding fashion. Here, the $\mathrm{P}_{\theta}(z|\hat{z_{i}})$ is zero, and the dynamic region prediction loss can be formulated as:
@@ 3.4 Inverse Dynamics via Denoising Diffusion Transformer
where $\varepsilon_{\theta}$ is the DiT denoiser, $\varepsilon\!\sim\!\mathcal{N}(0,I)$, $\bar{\alpha}_{\tau}$ follows a cosine noise schedule and $\mathbf{c}$ is the latent action embedding obtained from large language model. Inference is performed by drawing a Gaussian sample and running the learned reverse diffusion, yielding diverse yet physically plausible trajectories that close the perception–prediction–action loop.
@@ 4.1 Implementation Details
All models are implemented in PyTorch and trained on NVIDIA 8 A800 GPUs. We use an AdamW [114] optimizer with initial learning rate $10^{-3}$, weight decay $41e-$ 4, and a cosine learning-rate schedule with 5% linear warm-up. Batch size is set to 8, we set the query length of each modality $9$ and diffusion steps in DiT to 10. We weight the dynamic region, depth and segmentation prediction losses as $\lambda_{\text{dyn}}{=}0.1$, $\lambda_{\text{depth}}{=}0.001$, $\lambda_{\text{sem}}{=}0.1$, and the action loss as $\lambda_{\text{DiT}}{=}1$, respectively. We first pre-train DreamVLA on the language-free split of the CALVIN [115] and on the full DROID dataset [82]. The model predicts entire frames instead of comprehensive knowledge, keeping storage and computation requirements manageable. We then fine-tune DreamVLA on each target dataset using the comprehensive world knowledge forecasting objective. All models are trained for 20 epochs, and we select the checkpoint with the highest validation success rate (SR) for final evaluation.
@@ 4.2 Simulation Benchmark Experiments
Simulation setup. We evaluate DreamVLA on CALVIN [115], a simulated benchmark designed for learning long-horizon, language-conditioned robot manipulation policies. It comprises four distinct manipulation environments and over six hours of teleoperated play data per environment, captured from multiple sensors including static and gripper-mounted RGB-D cameras, tactile images, and proprioceptive readings. We report the success rate of every track and the average length of 5 tasks.
Results. As shown in Table 1, DreamVLA achieves the highest performance on ABC-D tasks, indicating that our method has better multi-task learning and generalization capabilities in simulation tasks. More detailed experimental results are shown in the supplementary materials.
@@ 4.4 Ablation Study
[…]
@@ A.1 DreamVLA Architecture
To retain spatial correspondence, we inject fixed sine–cosine positional encodings into the token embeddings. These tokens are then processed through several Transformer encoder layers, followed by a modality-specific linear projection head that maps each patch token to its output space, such as per-pixel depth values or semantic logits—thereby reconstructing the expected visual signals of future observations. Concrete details of each module are shown in Table  9.
@@ B.2 Simulation Results
In particular, our model significantly outperforms two-stage inverse dynamics approaches such as Susie [116], demonstrating the effectiveness of our end-to-end architecture that unifies multimodal prediction and action generation. Compared to CLOVER [118], UP-VLA [57], Seer [56], which also incorporates visual foresight, DreamVLA benefits from a more integrated design and joint optimization, resulting in consistently stronger execution accuracy. Furthermore, our method surpasses video generation-based pretraining approaches like GR-1 [14], highlighting the advantage of coupling visual world modeling with action planning in a single framework.

### v3

@@ 2.1 Vision–Language–Action Models
The earliest VLA [16, 71, 2, 72, 73, 74] lay the foundation by combining pretrained vision-language representations with task-conditioned policies for manipulation and control. Inspired by the recent advances of Large Language Models [75, 76, 77, 78] and multimodal large language models [28, 26, 79, 65, 80] and the emergence of large-scale robot datasets [12, 81, 82, 83], VLA has become a trend in robot learning. RT series [2, 84, 85] is the pioneer attempt to fine-tune the MLLM on robot demonstration datasets, resulting in strong accuracy and generalization. Building on this foundation, many advanced techniques [30, 32, 13, 1, 33, 34, 73, 35, 36, 37, 86, 87, 88, 38, 89] are developed to boost the performance. Meanwhile, considering the advantage of the diffusion model in modeling multi-peak, some researchers [90, 91, 92, 93, 94] employ different architectures to sample action from noise conditioned on observation, task instruction, and robot prior knowledge. Given on this manner which directly maps observation and instruction to action lacks reasoning steps like LLM [62], most existing methods [43, 5, 44, 45, 46, 47, 48, 49] leverage a copilot image/video generation model to generate future frames then predict action sequences conditioned on goal images. However, the above methods still need an extra generation model, which introduces inference time and computation load. Therefore, several methods [56, 57, 58, 59, 60, 61] integrate pixel-level forecasting with the action prediction in a single framework, which exploits the synergy of prediction and planning. Despite success, these methods naturally exhibit limitations in redundant reconstruction [95], and lack spatial and semantic information.
@@ 2.2 Knowledge Forecasting for Robotics
Learning future world knowledge for robot training has increasingly become popular to enable policies for achieving an action-forecasting loop. Early attempts [49, 19, 14, 43, 51, 50, 96] to implement this based on off-the-shelf video generation models [97, 53] and feed the goal images or states into policy model to conduct inverse dynamics. This two-stage training strategy is easy to implement but is limited by the performance and latency of video generation models. More advanced solutions couple forecasting with control by requiring the policy to produce, in addition to actions, explicit predictions. Concretely, these works ask the policy to output (i) high-level subtask/option sequences or language plans that decompose long-horizon goals [98, 99, 100], (ii)latent future embeddings/latent actions that compactly encode forthcoming motor intentions [88], (iii)whole sub-goal images or short visual rollouts that anticipate how the scene should evolve [56, 58], and (iv) object-centric signals (e.g., bounding boxes) that capture manipulation-relevant dynamics [83, 87]. This line of work demonstrates better performance and generalization. However, the future states are limited to redundant visual information [63, 64, 101, 69, 102, 66] or monotonous states [21, 48]. In contrast to previous work, DreamVLA proposes to predict future knowledge in an efficient (dynamic region) and effective (comprehensive knowledge) way, demonstrating strong performance and generalization.
@@ 3.1 Problem Definition and Notation
$\mathbf{w}_{t+n}=\mathcal{M}\left(l,o_{t},s_{t}|\text{{<dream>}}\right).$ (1)
@@
$\hat{a}_{t:t+n-1}=\mathcal{D}(\mathcal{M}\bigl{(}l,o_{t},s_{t},\texttt{<dream>}|\texttt{<action>})),$ (3)
@@ 3.3 Comprehensive World Knowledge Prediction
$),\displaystyle\sum_{(z_{i},\tilde{z}_{i})\in\mathcal{D}}\log\mathrm{P}(x_{i}|\tilde{x}_{i})\geq\sum_{(x_{i},\tilde{x}_{i})\in\mathcal{D}}\Big{(}\mathbb{E}_{z_{i}\sim\mathrm{Q}_{\phi}(\mathbf{z}|x_{i})}\big{[}\log\mathrm{P}_{\psi}(x_{i}|z_{i})\big{]}-D_{\mathrm{KL}}\big{[}z,\mathrm{P}_{\theta}(\mathbf{z}|\hat{z_{i}}$$\big{]}\Big{)}$, (4)
[…]
@@ 3.4 Inverse Dynamics via Denoising Diffusion Transformer
where $\varepsilon_{\theta}$ is the DiT denoiser, $\varepsilon\!\sim\!\mathcal{N}(0,I)$, $\bar{\alpha}_{\tau}$ follows a cosine noise schedule and $\mathbf{c}$ is the latent action embedding obtained from a large language model. Inference is performed by drawing a Gaussian sample and running the learned reverse diffusion, yielding diverse yet physically plausible trajectories that close the perception–prediction–action loop.
@@ 4.1 Implementation Details
All models are implemented in PyTorch and trained on NVIDIA 8 A800 GPUs. We use an AdamW [116] optimizer with initial learning rate $10^{-3}$, weight decay 1e41e-4, and a cosine learning-rate schedule with 5% linear warm-up. Batch size is set to 64, we set the query length of each modality $9$ and diffusion steps in DiT to 10. We weight the dynamic region, depth and segmentation prediction losses as $\lambda_{\text{dyn}}{=}0.1$, $\lambda_{\text{depth}}{=}0.001$, $\lambda_{\text{sem}}{=}0.1$, and the action loss as $\lambda_{\text{DiT}}{=}1$, respectively. We first pre-train DreamVLA on the language-free split of the CALVIN [117] and on the full DROID dataset [82]. For the LIBERO benchmark, we first pretrain DreamVLA on LIBERO-90 and then finetune on each track. The model predicts entire frames instead of comprehensive knowledge, keeping storage and computation requirements manageable. We then fine-tune DreamVLA on each target dataset using the comprehensive world knowledge forecasting objective. All models are trained for 20 epochs, and we select the checkpoint with the highest validation success rate (SR) for final evaluation.
@@ 4.2 Simulation Benchmark Experiments
Simulation setup. We evaluate DreamVLA on CALVIN [117] and LIBERO [122] benchmark. CALVIN is a simulated benchmark designed for learning long-horizon, language-conditioned robot manipulation policies. It comprises four distinct manipulation environments and over six hours of teleoperated play data per environment, captured from multiple sensors including static and gripper-mounted RGB-D cameras, tactile images, and proprioceptive readings. We report the success rate of every track and the average length of 5 tasks. Additionally, evaluations are also conducted on LIBERO [122], a simulated benchmark spanning four suites (LIBERO-Spatial/-Object/-Goal/-Long). Each suite contains 10 tasks supported by 50 human-teleoperated demonstrations, targeting spatial reasoning, object-centric manipulation, and goal completion.
Results. As shown in Table 1, DreamVLA achieves the highest performance on ABC-D tasks, Our method surpasses Roboflamingo [30], 3D Diffusor Actor [93], OpenVLA [1], RoboDual [119], UNIVLA [120], Robovlm [37] and GR1 [14], which directly projects the RGB/depth image to action signals as shown in Fig. 1(a) in the manuscripts. Compared to methods that use a copilot model to generate sub-goal images as input, like Susie [118] and CLOVER [121] as shown in Fig. 1(b) in manuscripts, our model significantly achieves more accurate control. DreamVLA outperforms approaches like UP-VLA [57], Seer [56], and VPP [49] as shown in Fig. 1(c) in manuscripts, which merge whole sub-goal image foresight into one VLA to take benefits from a more integrated design and joint optimization. indicating that our method has better multi-task learning and generalization capabilities in simulation tasks. For the LIBERO benchmark [122], DreamVLA exhibits better or comparable ability across all tracks compared to previous approaches by future world knowledge prediction as shown in Table 2.
Table 2: The extended LIBERO experiments. DreamVLA achieves the best or competitive performance across all tracks compared to previous approaches. The best results are bolded.
Methods Scores (%) Average
Spatial Object Goal Long
Diffusion Policy [90] 78.3 92.5 68.3 50.5 72.4
Octo [13] 78.9 85.7 84.6 51.1 75.1
OpenVLA [1] 84.7 88.4 79.2 53.7 76.5
SpatialVLA [36] 88.2 89.9 78.6 55.5 78.1
CoT-VLA [58] 81.1 87.5 91.6 87.6 69.0
DreamVLA 97.5 94.0 89.5 89.5 92.6
@@ 4.4 Ablation Study
Each <dream> query contains three groups of elements: dynamic, depth, and semantics, each assigned $K$ queries. We vary $K\in\{4,\,9,\,16\}$ to examine its influence. When $K=4$, the limited capacity prevents the model from encoding fine-grained motion, geometry, and semantics, so accuracy drops even though memory usage is lowest. With $K=9$, each modality has sufficient bandwidth without overloading the backbone, yielding the best success rate and the longest uninterrupted task execution. Increasing to $K=16$ introduces redundant tokens that compete for attention and raise GPU memory, bringing no extra gain and slightly lower generalization.
@@ A.1 DreamVLA Architecture
To retain spatial correspondence, we inject fixed sine–cosine positional encodings into the token embeddings. These tokens are then processed through several Transformer encoder layers, followed by a modality-specific linear projection head that maps each patch token to its output space, such as per-pixel depth values or semantic logits—thereby reconstructing the expected visual signals of future observations. Concrete details of each module are shown in Table  10.
@@ B.2 Simulation Results
In particular, our model significantly outperforms two-stage inverse dynamics approaches such as Susie [118], demonstrating the effectiveness of our end-to-end architecture that unifies multimodal prediction and action generation. Compared to CLOVER [121], UP-VLA [57], Seer [56], which also incorporates visual foresight, DreamVLA benefits from a more integrated design and joint optimization, resulting in consistently stronger execution accuracy. Furthermore, our method surpasses video generation-based pretraining approaches like GR-1 [14], highlighting the advantage of coupling vision prediction with action planning in a single framework.
@@ B.4 Real-world Settings
B.5 Inference latency
model part inference time
image, text and state encoders 12 ms
observation forward pass w/dream query 19 ms
w/o dream query 16 ms
action forward pass (10 step) 60 ms
total 91 ms
w/o dream query 88 ms
Table 13: Inference time of our model on a NVIDIA GeForce RTX 4090 GPU, we test 5 times and take average time.
Table 13 reports end-to-end latency for processing two camera images on an NVIDIA GeForce RTX 4090. At inference time, no explicit image decoding is required, and the system runs at 11 Hz. The results show: (i) Auxiliary cues incur minimal overhead. Our “dream queries” are token-level predictions (no explicit image decoding and no external models). The incremental cost is 3 ms (3.4%), i.e., 91 ms vs. 88 ms without dream queries. (ii) Latency is dominated by the action head rather than the auxiliary cues. A 10-step action head contributes about 60 ms. This cost scales with the number of sampling steps and model size; it can be reduced by using fewer steps, a smaller DiT variant, faster samplers. (iii) For latency-critical applications, we can prune dream queries (e.g., keep only dynamic regions, or dynamic+depth) and/or increase action chunking or run the action head asynchronously to amortize computation, without changing the observation pathway.